# RHINO: A large horn antenna for detecting the 21cm global signalRHINO: A large horn antenna for detecting the 21cm global signal

### v1

@@ 2.3 Electromagnetic simulations
Following this process, we arrived at the reference design shown in Fig. 1. Antenna patterns are shown in Figs. 2, 3, and 4. The horn aperture is $7.3\times 6.0$m, with a flare length (along the corner of the flare section) of 5.1m, for a flare section vertical height of 4.3m. The waveguide has an aperture of $2.0\times 3.5$m and a vertical height of 3m. This has a gain of 12.5 dB and FWHM of $44^{\circ}$ at 65 MHz (Figs. 2 and 4), and achieves a return loss of $S_{11}<-10$ dB across the design band (Fig. 5). The (relative) backlobe level at 65 MHz is $-18.4$ dB however, and the sidelobe level is around $-15$ dB, both of which are outside the design requirements. This behaviour is expected; in a rectangular horn antenna, the electric field distribution over the E-plane walls is uniform, and as a result, large sidelobe levels in the far-field radiation patterns of the corresponding planes will be observed (Kraus & Marhefka, 2002) unless the E-field distribution can be modified.
@@
To recap, the aim of this design exercise was to find the simplest horn antenna that could approach the requirements above. While a simple and workable design was found, it falls outside the target requirements for the sidelobe and backlobe levels, and so in practice a refined design is needed. Several elaborations are possible, and will be explored through detailed simulations in a future work. Briefly, these include the addition of choke rings at the aperture (to suppress the backlobe), and a variable flare angle and/or corrugations around the transition from the flare section to the aperture, to improve matching, and hence reduce sidelobes and improve bandwidth. A densely corrugated design is likely to add significant additional complexity and expense for a horn approaching 10m in size, and so we do not consider this option, although we note that corrugated designs have been demonstrated for horns of a couple of metres in size used by the BINGO experiment (Wuensche et al., 2022). Other options include the construction of an extensive ground screen, to suppress emission from low elevations and below the horizon, as well as more sophisticated optics such as lenses.
@@ 2.4 Mechanical design
The conductivity of the surface is also important, as Ohmic losses will reduce the antenna efficiency. Findlay (1990) used welded sheet metal to minimise the losses in a pyramidal horn antenna with aperture $5.4\times 4.0$m and slant length 36.7m for example. For a large vertical antenna, however, this presents challenges in terms of the necessary support structure, as well as wind loading, thermal expansion etc. An alternative is to use welded mesh, which has benefits in terms of reduced cost and weight, but will have higher losses. It is also less straightforward to achieve suitable conductive joins between pieces of mesh, as the narrow wire is relatively fragile for welding. These issues are also being studied using the scaled-down prototype, which uses 13 mm galvanised steel mesh (0.89 mm wire gauge).
For the support structure, simple options such as scaffolding and telegraph poles have been considered, which require minimal foundation work. The prototype uses wooden beams to shape the horn, with mesh stapled onto the inner surface, and metal scaffolding poles to support the structure. For the full-sized horn, an aluminium frame made up of several sections will be used instead.
@@
Figure 7: Slices through the normalised antenna pattern as in Fig. 2, but now for the corrugated conical horn. The sidelobe level and degree of symmetry are significantly improved in this design.
@@ 2.5 Alternative corrugated conical horn design
Another option is to design a corrugated horn, which provides many advantages compared to a conventional pyramidal horn, at the cost of a more complicated design. Fig. 6 shows a proposed alternative horn geometry, consisting of a conical horn with a small number of large corrugations (five in this iteration). This should be contrasted with more conventional corrugated designs at higher frequencies, which place many corrugations over along the flare section of the horn. Some rules of thumb for corrugation depth and spacing are discussed in Lawrie & Peters (1966) for instance.
These corrugations are nevertheless successful in significantly improving the antenna pattern, to the extent that it now satisfies all of the design requirements in Sect. 2.2. Slices through the antenna patterns are shown in Fig. 7. The most noticeable improvements are in terms of the sidelobe level, which is now better than $-30$ dB in both the E-plane and H-plane, with a high degree of symmetry. The backlobe is also greatly reduced, and is better than $-30$ dB across the frequency range.
@@ 3.3 Continuous wave source
One of the novel aspects of our calibration scheme is the inclusion of a continuous wave (CW) source. This injects a constant tone into the system that has a large amplitude but only occupies a single frequency channel, leaving most of the band unaffected, and barely increasing the total power into the amplifier. The CW source is monitored in real-time by a logarithmic power meter to track its stability, and can be tuned to different frequencies on demand. For the prototype system, we use an Analog Devices ADF4351 (square wave) synthesiser board attached to an AD8317 power meter via a splitter. The CW module is then connected to the receiver system via a combiner, which sums the CW and antenna signals.
@@
This also relies on the bandpass being stable on relatively long timescales, i.e. the correlated gain fluctuations are assumed to be strongly correlated in frequency (Pollak et al., 2019). This assumption can be tested by changing the frequency of the CW source, or even injecting a comb of multiple CW signals across the band. Our expectation is that this assumption should be good enough to significantly suppress the correlated gain fluctuations on intermediate timescales however. The calibration loads will still be switched in periodically to re-measure the bandpass and re-calibrate the CW source amplitude, but this is intended to be done much less often than in systems that do not have CW sources.
@@ 3.4 Filters and low-noise amplifiers
RHINO targets a smaller fractional bandwidth than many other 21cm global signal experiments, in part due to the limitations of the basic horn design, and in part due to the anticipated RFI environment. The design band is 60 – 85 MHz (34% fractional bandwidth), with a steep rolloff required at the top of the band to suppress bright FM radio transmission that begins at 87.5 MHz. A Mini-Circuits SBP-70+ bandpass filter satisfies our requirements. In addition, a Nooelec Flamingo FM bandstop filter [^4] https://www.nooelec.com/store/downloads/dl/file/id/95/product/291/flamingo_fm_filter_datasheet_revision_1.pdf is included, which has an attenuation of $-30$ dB or better across the FM band.
We use Mini-Circuits ZKL-2+ low-noise amplifiers [^5] https://www.minicircuits.com/pdfs/ZKL-2.pdf (LNA), with a typical gain of 31.1 dB, an input VSWR of 1.02:1, an output VSWR of 1.19:1, and a noise figure of 3.4 dB, all at a reference frequency of 70 MHz. Two of these are chained together with 3 dB attenuators attached to the input and output in order to reduce reflections (the first one uses the power combiner to provide this level of attenuation at the input). This model of LNA is packaged in a metal case as standard, providing excellent shielding. We attach this to an aluminium plate to help stabilise its temperature, but do not directly control the LNA temperature beyond heating the inside of the receiver box to around $40^{\circ}$ to isolate it from fluctuations in the outside temperature.
@@ 3.5 Integrated vector network analyser
A vector network analyser (VNA) is integrated into the system so that reflection coefficients along several signal paths can be measured in situ, rather than relying on lab measurements that occur under different environmental conditions. At the moment, this is a NanoVNA-H4666https://nanovna.com/?page_id=21 with a USB control interface. The VNA is connected to a set of SOLT (short-open-load-through) calibration targets via the main 8-way switch and an associated 2-way switch so that it can be be periodically re-calibrated in-situ.
Depending on the settings of the switches, the VNA can be used to probe reflection coefficients along paths to the receiver, the antenna, and the calibration loads.
@@ 4.1 Continuous wave calibration scheme
The intention is for the relevant reflection coefficients to be measured in situ by periodically switching in the built-in VNA (e.g. de Lera Acedo et al., 2022).
@@
Here, $g$ is the gain due to the filters and amplifiers within the receiver, $T_{\rm n}$ is the noise added by the receiver, and the antenna temperature and noise wave components are given by (Meys, 1978; Monsalve et al., 2017a; Sun et al., 2024)
@@
This is independent of frequency. Note that schemes based on differencing can be biased by noise fluctuations and smooth (but non-zero) variations in the continuum power between neighbouring channels. Instead of directly differencing with neighbouring channels, one can instead fit a continuum across multiple channels and de-trend. Using multiple channels also helps average down effects due to noise. Beyond this, a full forward-modelling approach to the calibration permits self-consistent estimates of these effects to be made (c.f. Roque et al., 2021). The real-world performance of the RHINO CW calibration, and different algorithmic approaches to measuring correlated gain fluctuations, will be examined in a future work.
The CW approach can track temporal variations in the gains that are strongly correlated across the band, but would become ineffective if there is substantial spectral variation in the $1/f$ noise. Related issues were examined by Pollak et al. (2019), and solutions such as injecting a frequency comb (instead of a single tone) can be explored if necessary. Establishing whether the gain fluctuations are indeed strongly correlated in frequency is a key priority for the RHINO prototype phase.
@@ 4.2 Warm calibration loads
A bandpass calibration is also needed, and cannot be provided by the CW source. For this, we use the more traditional approach of periodically observing a termination heated to a known temperature, i.e. a warm or hot calibration load. This will produce a total noise power of $P_{\rm load}=k_{B}T_{\rm load}\Delta\nu$ over a bandwidth $\Delta\nu$. Because we do not need to rapidly switch between the calibration load and the antenna, we can switch to the load periodically, dwell on it for longer, and get a lower-noise measurement of its reference power. If the effective temperature of the load is well-measured, and any additional spectral shape due to reflections is known, this gives us a bandpass calibration. With two loads of different temperatures to switch between, we can derive an absolute temperature calibration.
@@
where $T_{\rm calsrc}$ is the effective temperature of the termination. This is derived from the physical temperature of the termination plus a correction due to the temperature of the (lossy) cable that connects it to the receiver. For a linear temperature gradient between the calibration source and the receiver’s input port, and a cable insertion loss below 0.5 dB, we can make the approximation (Findlay, 1966),

### v2

@@ 2.3 Electromagnetic simulations
Following this process, we arrived at the reference design shown in Fig. 1. Antenna patterns are shown in Figs. 2, 3, and 4. The horn aperture is $7.3\times 6.0$m, with a flare length (along the corner of the flare section) of 5.1m, for a flare section vertical height of 4.3m. The waveguide has an aperture of $2.0\times 3.5$m and a vertical height of 3m. This has a gain of 12.5 dB and FWHM of $44^{\circ}$ at 65 MHz (Figs. 2 and 4), and achieves a return loss of $S_{11}<-10$ dB across the design band (Fig. 5). The (relative) backlobe level at 65 MHz is 18.4-18.4 dB however, and the sidelobe level is around 15-15 dB, both of which are outside the design requirements. This behaviour is expected; in a rectangular horn antenna, the electric field distribution over the E-plane walls is uniform, and as a result, large sidelobe levels in the far-field radiation patterns of the corresponding planes will be observed (kraus2002antennas) unless the E-field distribution can be modified.
@@
To recap, the aim of this design exercise was to find the simplest horn antenna that could approach the requirements above. While a simple and workable design was found, it falls outside the target requirements for the sidelobe and backlobe levels, and so in practice a refined design is needed. Several elaborations are possible, and will be explored through detailed simulations in a future work. Briefly, these include the addition of choke rings at the aperture (to suppress the backlobe), and a variable flare angle and/or corrugations around the transition from the flare section to the aperture, to improve matching, and hence reduce sidelobes and improve bandwidth. A densely corrugated design is likely to add significant additional complexity and expense for a horn approaching 10m in size (see Sect. 2.6), and so we have not adopted this option, although we note that corrugated designs have been demonstrated for horns of a couple of metres in size used by the BINGO experiment (Wuensche:2021dcx). Other options include the construction of an extensive ground screen, to suppress emission from low elevations and below the horizon, as well as more sophisticated optics such as lenses.
@@
2.4 Sensitivity to ground conditions
One of the main motivations for using a horn antenna instead of a compact antenna (such as a dipole) is that horns are expected to be less sensitive to the detailed conditions of the ground or ground plane, making it less important to model these components with high accuracy. In this section, we present illustrative EM simulations to examine this expectation, with a more detailed investigation to be presented in Elmakadema et al. (in prep.).
Figure 6: Beam correction factor as a function of frequency, for a single reference time, for three different assumptions about the ground conditions beneath the horn antenna: vacuum (no ground); uniform dry soil ($\rho=0.001$ S m-1, $\epsilon_{r}=2$); and uniform wet soil ($\rho=0.1$ S m-1, $\epsilon_{r}=7$).
To study the sensitivity to ground conditions, we use the same horn antenna design as in Sect. 2.2, but alter the boundary conditions to represent three different idealised ground scenarios – vacuum (the same as the models shown in the previous section), wet soil, and dry soil. For the non-vacuum cases, the EM simulation domain is a hemisphere of radius 20m immediately beneath the bottom edge of the waveguide, with a uniform conductivity and dielectric constant representative of either wet or dry soil, as reported in 2022MNRAS.515.1580S. These conditions are intended to bracket reasonable wet and dry conditions, with ‘typical’ soil conditions falling in between depending on the site. Convergence tests were performed to ensure that the choice of domain and mesh parameters in the simulation are appropriate (c.f. Mahesh:2021rly).
Next, we calculate the beam correction factor for each scenario, using the reprocessed Haslam 408 MHz map (Haslam1982; 2015MNRAS.451.4311R) scaled by a synchrotron power-law spectrum as our sky brightness temperature model. The beam correction factor is calculated as defined in 2022MNRAS.515.1580S,
$B_{\rm corr}(\nu,t)=\frac{\int_{\Omega}T_{\rm sky}(\nu_{\rm ref},\hat{n})B(\nu,t,\hat{n})d\Omega}{\int_{\Omega}T_{\rm sky}(\nu_{\rm ref},\hat{n})B(\nu_{\rm ref},t,\hat{n})d\Omega}\frac{\int_{\Omega}B(\nu_{\rm ref},\hat{n})d\Omega}{\int_{\Omega}B(\nu,\hat{n})d\Omega}.$ (4)
Here, $B(\nu,t)$ is the power beam pattern as a function of frequency and local sidereal time of observation, $t$, and $T_{\rm sky}$ is the sky brightness model. A reference frequency of $\nu_{\rm ref}=70$ MHz has been chosen in what follows. This factor accounts for the frequency dependence of the mapping between antenna temperature and brightness temperature, and the application of this correction is the main part of a 21cm global signal analysis where a precise beam model is needed. The correction factor is relatively spectrally smooth for dipole-type antennas (e.g. see 2021MNRAS.506.2041A), which have antenna patterns dominated by a mainlobe that is relatively stable with frequency. Interactions between the antenna and its ground plane or the ground itself (for example) can contribute additional chromaticity however (2019ApJ...874..153B; 2022MNRAS.515.1580S). Horns and other aperture antennas, on the other hand, are intrinsically more chromatic, and tend to have more complex sidelobes, but do not conventionally require a ground plane to be incorporated in the design. Most of their sensitivity to the ground comes through far sidelobes caused by diffraction around the aperture, which can be mitigated through design choices to minimise sidelobes for example.
Fig. 6 shows the beam correction factor for the three scenarios at two local sidereal times: LST = 20h, when a cold region of the sky is at zenith, and LST = 08h, when the Galactic plane is transiting. Focusing first on the 20h case, we see that the beam correction factor changes relatively little between the wet and dry cases, which supports the expectation that the horn antenna design should be insensitive to ground conditions. There is a larger difference between these cases and the vacuum case, as the change in boundary conditions now allows for reflections from the ground that are not possible in vacuum. For the 08h case, when the bright Galactic plane is overhead, the differences between the three cases become more pronounced, but are still at the 1% level or less. This is comparable to the results of 2022MNRAS.515.1580S (see their Fig. 11), which shows beam correction factor variations of around $\pm 0.3\%$ between the wet and dry cases for different shapes and sizes of ground plane (in this case, for a LEDA triangular dipole antenna).
@@ 2.5 Mechanical design
The conductivity of the surface is also important, as Ohmic losses will reduce the antenna efficiency. Findlay1990 used welded sheet metal to minimise the losses in a pyramidal horn antenna with aperture $5.4\times 4.0$m and slant length 36.7m for example. For a large vertical antenna, however, this presents challenges in terms of the necessary support structure, as well as wind loading, thermal expansion etc. An alternative is to use welded mesh, which has benefits in terms of reduced cost and weight, but will have higher losses. It is also less straightforward to achieve suitable conductive joins between pieces of mesh, as the narrow wire is relatively fragile for welding. To overcome this, we selected 8g (4mm diameter) mild steel welded mesh panels with $50\times 50$ mm mesh aperture. Each $2.4\times 1.2$ m mesh panel is cut to size and then welded into a box frame made from $50\times 50\times 5$ mm steel angle. The mesh gauge is sufficiently thick to permit efficient direct welds onto the angle. The panels can then be bolted together so that the flat angle surfaces perpendicular to the mesh are directly in contact. Fig. 1 shows a detailed design for this arrangement, consisting of 59 panels for a total mass of around 1800 kg. This structure is self-supporting, but requires bracing with guy lines or similar to withstand wind loading etc.
@@ 2.6 Alternative corrugated conical horn design
Another option is to design a corrugated horn, which provides many advantages compared to a conventional pyramidal horn, at the cost of a more complicated design. We present a suitably optimised alternative design here to illustrate the differences in performance and complexity of construction, and to provide a reference future development path for an enhanced low-frequency absolutely-calibrated horn antenna experiment. Fig. 7 shows an alternative horn geometry, consisting of a conical horn with a small number of large corrugations (five in this iteration). This should be contrasted with more conventional corrugated designs at higher frequencies, which place many corrugations over along the flare section of the horn. Some rules of thumb for corrugation depth and spacing are discussed in 1966ITAP...14..605L for instance.
Figure 8: Slices through the normalised antenna pattern as in Fig. 2, but now for the corrugated conical horn shown in Fig. 7. The sidelobe level and degree of symmetry are significantly improved in this design.
These corrugations are nevertheless successful in significantly improving the antenna pattern, to the extent that it now satisfies all of the design requirements in Sect. 2.2. Slices through the antenna patterns are shown in Fig. 8. The most noticeable improvements are in terms of the sidelobe level, which is now better than 30-30 dB in both the E-plane and H-plane, with a high degree of symmetry. The backlobe is also greatly reduced, and is better than 30-30 dB across the frequency range.
@@ 3.3 Continuous wave source
One of the novel aspects of our calibration scheme (as far as ground-based 21cm global signal experiments are concerned) [^4] Continuous wave or ‘tone’ calibration methods have previously been used in a number of related contexts (e.g. bale2016fields; 2021arXiv210305085B). is the inclusion of a continuous wave (CW) source. This injects a constant tone into the system that has a large amplitude but only occupies a single frequency channel, leaving most of the band unaffected, and barely increasing the total power into the amplifier. The CW source is monitored in real-time by a logarithmic power meter to track its stability, and can be tuned to different frequencies on demand. For the prototype system, we use an Analog Devices ADF4351 (square wave) synthesiser board attached to an AD8317 power meter via a splitter. The CW module is then connected to the receiver system via a combiner, which sums the CW and antenna signals.
@@
This also relies on the bandpass being stable on relatively long timescales, i.e. the correlated gain fluctuations are assumed to be strongly correlated in frequency (2019MNRAS.489..548P). This assumption can be tested by changing the frequency of the CW source, or even injecting a comb of multiple CW signals across the band. Our expectation is that this assumption should be good enough to significantly suppress the correlated gain fluctuations on intermediate timescales however. The calibration loads will still be switched in periodically to re-measure the bandpass and re-calibrate the CW source amplitude, but this is intended to be done much less often than in systems that do not have CW sources.
@@ 3.4 Filters and low-noise amplifiers
RHINO targets a smaller fractional bandwidth than many other 21cm global signal experiments, in part due to the limitations of the basic horn design, and in part due to the anticipated RFI environment. The design band is 60 – 85 MHz (34% fractional bandwidth), with a steep rolloff required at the top of the band to suppress bright FM radio transmission that begins at 87.5 MHz. A Mini-Circuits SBP-70+ bandpass filter satisfies our requirements. In addition, a Nooelec Flamingo FM bandstop filter [^5] https://www.nooelec.com/store/downloads/dl/file/id/95/product/291/flamingo_fm_filter_datasheet_revision_1.pdf is included, which has an attenuation of 30-30 dB or better across the FM band.
We use Mini-Circuits ZKL-2+ low-noise amplifiers [^6] https://www.minicircuits.com/pdfs/ZKL-2.pdf (LNA), with a typical gain of 31.1 dB, an input VSWR of 1.02:1, an output VSWR of 1.19:1, and a noise figure of 3.4 dB, all at a reference frequency of 70 MHz. Two of these are chained together with 3 dB attenuators attached to the input and output in order to reduce reflections (the first one uses the power combiner to provide this level of attenuation at the input). This model of LNA is packaged in a metal case as standard, providing excellent shielding. We attach this to an aluminium plate to help stabilise its temperature, but do not directly control the LNA temperature beyond heating the inside of the receiver box to around $40^{\circ}$C to isolate it from fluctuations in the outside temperature.
@@ 3.5 Integrated vector network analyser
A vector network analyser (VNA) is integrated into the system so that reflection coefficients along several signal paths can be measured in situ, rather than relying on lab measurements that occur under different environmental conditions. At the moment, this is a NanoVNA-H4777https://nanovna.com/?page_id=21 with a USB control interface. The VNA is connected to a set of SOLT (short-open-load-through) calibration targets via the main 8-way switch and an associated 2-way switch so that it can be periodically re-calibrated in-situ.
Depending on the settings of the switches, the VNA can be used to probe reflection coefficients along paths to the receiver, the antenna, and the calibration loads. An important consideration is where the reference plane of the VNA is defined. Laboratory measurements of the switches and cables may be needed to permit the VNA measurements to be transferred to a different reference plane (e.g. 2025ExA....59....7R). In turn, this may require enhanced measures to stabilise the switch and cable properties (e.g. see Zerafa) to prevent the transfer coefficients from shifting undetected.
@@ 4.1 Continuous wave calibration scheme
The intention is for the relevant reflection coefficients to be measured in situ by periodically switching in the built-in VNA (e.g. 2022NatAs...6..984D).
@@
Here, $g$ is the gain due to the filters and amplifiers within the receiver, $T_{\rm n}$ is the noise added by the receiver, and the antenna temperature and noise wave components are given by (1978ITMTT..26...34M; 2017ApJ...835...49M; 2024Univ...10..236S)
@@
This is independent of frequency. Note that schemes based on differencing can be biased by noise fluctuations and smooth (but non-zero) variations in the continuum power between neighbouring channels. Instead of directly differencing with neighbouring channels, one can instead fit a continuum across multiple channels and de-trend. Using multiple channels also helps average down effects due to noise. Beyond this, a full forward-modelling approach to the calibration permits self-consistent estimates of these effects to be made (c.f. 2021MNRAS.505.2638R). The real-world performance of the RHINO CW calibration, and different algorithmic approaches to measuring correlated gain fluctuations, will be examined in a future work.
The CW approach can track temporal variations in the gains that are strongly correlated across the band, but would become ineffective if there is substantial spectral variation in the $1/f$ noise. Related issues were examined by 2019MNRAS.489..548P, and solutions such as injecting a frequency comb (instead of a single tone) can be explored if necessary. Establishing whether the gain fluctuations are indeed strongly correlated in frequency is a key priority for the RHINO prototype phase.
@@ 4.2 Warm calibration loads
A bandpass calibration is also needed, and cannot be provided by the CW source. For this, we use the more traditional approach of periodically observing a termination heated to a known temperature, i.e. a warm or hot calibration load. This will produce a total noise power of $P_{\rm load}=k_{\rm B}T_{\rm load}\Delta\nu$ over a bandwidth $\Delta\nu$. Because we do not need to rapidly switch between the calibration load and the antenna, we can switch to the load periodically, dwell on it for longer, and get a lower-noise measurement of its reference power. If the effective temperature of the load is well-measured, and any additional spectral shape due to reflections is known, this gives us a bandpass calibration. With two loads of different temperatures to switch between, we can derive an absolute temperature calibration.
@@
where $T_{\rm calsrc}$ is the effective temperature of the termination. This is derived from the physical temperature of the termination plus a correction due to the temperature of the (lossy) cable that connects it to the receiver. For a linear temperature gradient between the calibration source and the receiver’s input port, and a cable insertion loss below 0.5 dB, we can make the approximation (1966ARA&A...4...77F),